# New black hole mergers in the LIGO–Virgo O3 data from a gravitational wave search including higher-order harmonics

### v1

@@ III.3 Comparison to the LVK and OGC -only catalogs
In Table 3, we compare our pipeline’s results for the O3 Hanford–Livingston coincident events published in the GWTC-3 catalog333https://gwosc.org/GWTC-3/ Abbott et al. (2021). The GWTC-3 catalog combines results from four pipelines: three that search using templates which only contain the mode: GstLAL, Multi-Band Template Analysis (MBTA) and PyCBC Sachdev et al. (2019); Usman et al. (2016); Andres et al. (2022); and one that searches for transient signals with minimal assumptions about sources, cWB Klimenko et al. (2016). We also include the comparison with the 4-OGC catalog taken from the summary data in the repository444https://github.com/gwastro/4-ogc/blob/master/search/4OGC_top.txt listed in Nitz et al. (2021). Note that the quoted GWTC-3 squared SNR, , is from LVK’s PE runs for the events. The search SNR is generally lower, and varies between the different pipelines that the LVK uses.
@@
We broadly recover the confident LVK detections, except some vetoed/missed events which are discussed in more detail in the subsections below. Overall, 9 events GWTC-3 events in Table 1 were vetoed in our pipeline. 4 GWTC-3 events fell below our trigger collection SNR threshold (GW190917_114630, GW191113_071753, GW200220_061928, GW200308_173609), note that these were confidently detected by only one of the LVK pipelines. Three GWTC-3 events were detected with in our pipeline (GW200208_222617, GW200306_093714, GW200322_091133), note again that these were confidently detected by only one of the LVK pipelines.
@@
At the end of Table 3, we include comparison with the list of 7 events in 4-OGC which are not present in GWTC-3 catalog (from which we recover 4 with and 2 were below our collection SNR threshold). GW190909_114149 is a marginal candidate in GWTC-3 which has from our pipeline. We still include it in Table 3 as it was first a significant candidate in GWTC-2 Abbott et al. (2021), but later updated to be in the marginal list The LIGO Scientific Collaboration et al. (2021). We also checked the overlap of our new candidate event list with the sub-threshold candidate list (i.e., which have but FAR day) released by the LVK collaboration: GWTC-2.1 and GWTC-3 The LIGO Scientific Collaboration et al. (2021); Abbott et al. (2021); we show the corresponding comparison in Table 5 in the Appendix. In the same table, we also show the overlap of our new candidate list with sub-threshold candidates from the IAS -only pipeline Olsen et al. (2022); Mehta et al. (2023).
III.3.1 Vetoed events
Our pipeline includes a series of signal consistency checks in order to veto triggers that are likely to arise from noise transients (glitches) to improve our search sensitivity (see section I of Venumadhav et al. (2019)). The vetoes are designed in a way that they are falsely triggered by real signals with only a low probability (a few percent) if their waveforms are accurately modeled by templates in the bank, and the noise is Gaussian. If a loud signal displays effects which are not present in our templates, like precession or eccentricity, the chances of the veto mis-identifying it as a glitch can be much higher. GW190521_030229, GW191109_010717 and GW200129_065458 were reported to be fairly high SNR with imprints of precession Abbott et al. (2020, 2021), which could be the reason for them being vetoed. For example, the presence of higher modes in GW190521 is disfavored with a Bayes factor of , while a precessing signal is favored with a factor of Abbott et al. (2020).
GW190701_203306 was vetoed due to excess power present after subtracting the best-fit waveform (note that the data around this event was treated for scattered light glitches Abbott et al. (2021)). Out of the veto tests that we apply to events, we noticed GW190828_065509 and GW190930_133541 each failed just one test (namely, the split test with chunk sets, and the test with 20 chunks, respectively).
We noticed a small issue in the pipeline where we artificially veto triggers very close ( s) to the edge of the timeseries files; this led to the event GW190725_174728 being vetoed. We will fix this issue in future searches. The event GW190426_190642 was detected confidently by only one of the LVK search pipelines and got vetoed in our case.
We expect to recover some of loud vetoed events if we also include precession in our template banks. Alternatively, one could formulate a procedure to override vetos or changing their thresholds based on injections with precessing waveforms. We leave exploring these directions to future work.
III.3.2 Other missed/downweighted events
@@
As we discussed earlier in Section II.2, we set a particular SNR threshold () to collect triggers from individual detectors. We do this in order to regulate the number of single-detector triggers that are passed to the more costly coincidence step of the pipeline. If we set the threshold based on the quadrature sum of the SNRs of the (orthogonalized) harmonics, we collect more triggers in a HM search as compared to a -only search with the same threshold. Hence, we raised the collection threshold (see Section II.2.1 for a detailed discussion). Due to this, we missed low-SNR events that were reported by -only pipelines in the previous literature. We label the events which fell below this threshold as “Below ” in the table, where the threshold depends on the number of templates in the bank and is therefore different for different banks. On further investigation, using a smaller value of is feasible with our current search setup, and thus we plan to lower the threshold in future searches.
@@
We do not include the following events in Table 3 because they were detected in Livingston–Virgo or Hanford–Virgo coincidence, or single detector search, all of which we have yet to run: GW190620_030421, GW190630_185205, GW190708_232457, GW190814, GW190910_112807, and GW190925_232845. We also only searched for BBH mergers in this paper, and did not recover the NSBH candidates GW191219_163120 and GW200115_042309 (in addition, the posterior probability for GW200210_092254 peaks in the NSBH region Abbott et al. (2021) and thus it was vetoed in our search). We hope to perform coincident searches which would also include Virgo (and KAGRA) in the future. We also intend to perform a separate search for NSBH mergers including our HM pipeline in a future paper.
GW191222_033537 was detected with in our pipeline but has comparatively a much higher false alarm rate. We checked that it was downweighted due to being in the proximity of a glitch (which led to the event acquiring unphysical amounts of SNR in HM).
@@ IV Interesting astrophysical properties of the new candidate events
In this section, we discuss interesting astrophysical properties of the 14 new candidate events, whose parameters we have reported in Table 1 and in Figs. 1, 2 and 3.
@@ IV.1 Masses in the IMBH range and pair instability mass gap
Black holes roughly within the mass range – are designated as intermediate-mass black holes (IMBHs). Several IMBH candidates have been suggested by electromagnetic observations, but these lack conclusive confirmation Greene et al. (2020). Gravitational waves have the potential to definitively detect objects in this regime, e.g., Abbott et al. (2020); Fragione et al. (2018). In our search, at % credibility, 4 candidate events have total masses in the IMBH range (GW190605_025957, GW190604_103812, GW191228_195619 and GW190530_133833). Among them, GW191228_195619 has the highest total mass (and also has the highest primary mass ), which makes it, by a large margin, the most massive candidate with detected in gravitational wave data so far.
Another interesting regime in BH masses is the pair instability mass gap. This is the regime in which stars undergo pair-instability supernova (where they are completely disrupted and no remnant is left behind). Otherwise, the stars undergo pulsational pair-instability supernova and iteratively lose their mass before collapsing to a black hole. These processes are expected to produce a gap roughly between – in the distribution of BH mass. Note however that the exact boundaries of such a gap are not well determined, as they depend on multiple uncertain factors such as the rate of carbon to oxygen burning in the core , angular momentum transport, metallicity, wind loss, etc. Marchant and Moriya (2020); Farmer et al. (2019); Mehta et al. (2022); Hendriks et al. (2023).
Recent analyses have not found any conclusive evidence for the upper mass gap in the observed BH mass distribution, e.g., Abbott et al. (2023); Roulet et al. (2021). This either challenges the current theoretical models of stellar evolution, or the gap might be filled by black holes formed via alternative mechanisms such as hierarchical mergers Anagnostou et al. (2020); Fragione et al. (2020); Veske et al. (2021); Fragione et al. (2021), or from the gaseous disk channel Tagawa et al. (2020); Samsing et al. (2022). In our search, at 90% CL, 9 candidates have median mass of the primary BH lying between – (# 1, 4, 5, 6, 8, 10, 11, 13, 14 in Table 1). One of our candidates: GW191228_195619 has its secondary BH mass overlapping with the mass-gap range. Our new candidate events could be helpful in estimating the abundance of BH (or rate of BH mergers) in/beyond the pair-instability mass regime. This can also shed light on BH progenitor channels.
@@ IV.2 High redshifts
In our search, we find that 9 of our new candidates have redshift at % CL (# 1, 2, 3, 6, 8, 10, 11, 13, 14 in Table 1). Finding high-redshift BHs can be important to answer questions such as: How does the merger rate of BHs evolve with redshift Fishbach et al. (2018)? Does the merger rate change in a different manner compared to the the star formation rate (which peaks at Madau and Dickinson (2014))? Furthermore, different BH formation channels can have different predictions for the merger rate at high redshift Mapelli et al. (2022); van Son et al. (2022). For example, the chemically homogeneous evolution channel produces BBHs of relatively higher masses and lower redshifts than the common envelope evolution channel Mandel and de Mink (2016). In the dynamical channel, massive BHs can have relatively longer delay times compared to star formation Mapelli et al. (2019). The redshifts of binaries can be also be correlated with other properties of BHs, e.g., spins Biscoveanu et al. (2022); Kushnir et al. (2016); Zaldarriaga et al. (2018). We aim to test some of these models in a future population inference study.
IV.3 Preference for positive effective spins
Measuring spins of the black holes is one of the best ways to distinguish between isolated and dynamical formation channels. In the isolated channel, accretion can lead to spins of BHs which are aligned in the direction of the binary angular momentum (leading to positive values of the effective spin parameter ). Dynamical channels, on the other hand, predict the astrophysical distribution to be symmetric around 0. This is because the the spins BH captured into a binary are not expected to be preferentially aligned along any particular direction. From the PE runs under our fiducial prior, 6 candidates have effective spin at credibility (# 2, 4, 6, 8, 9, 11 in Table 1). Note however that posteriors are significantly dependent on the prior as seen in the comparison in Fig. 3. Though, even under the isotropic spin prior, some of the new candidate events still have mild preference for .
Note that the preference for in our detection catalog does not immediately mean that astrophysics prefers this scenario, as this could be caused by selection effects. One of the selection effects is that positive binaries are louder than their negative counterparts; due to the orbital hangup effect, the positive binaries merge later and have power at higher frequencies Ng et al. (2018); Roulet and Zaldarriaga (2019). This effect is much more prominent for high mass binaries (where a significant portion of signal is masked by the low frequency end of the PSD) as we show in Fig. 6. There is another effect which further downweights high-mass negative signals. Their waveforms have a short duration in band and therefore can be more effectively mimicked by glitches. This degrades the sensitivity to such systems in our search Wadekar et al. (2023b) when we perform the background dependent reweighting in our ranking statistic (see Section II.3.3 and Ref. Wadekar et al. (2023b)). To robustly determine if the true astrophysical distribution also favors at high masses, we plan to a do a population study (where we will use injections with both positive and negative values to determine our pipeline selection function).
@@ IV.4 Asymmetric mass ratio systems
Five of our new candidates have median mass ratio (# 4, 6, 7, 10, 11 in Table 1). Finding sources with asymmetric mass ratios can shed light on the formation channels Abbott et al. (2020b). The globular cluster (dynamical) channel predicts that most merging BBHs have nearly equal masses Rodriguez et al. (2016). This is because higher mass BHs in a globular cluster sink to their cores due to dynamical friction and vice versa for lighter BHs. The caveat is that if clusters can retain merger products, they can produce “second-generation” mergers whose mass ratios can deviate significantly from unity Rodriguez et al. (2019). In the isolated binary channel, in certain cases, an initially asymmetric mass ratio system can evolve toward a more symmetric configuration due to mass accretion episodes Ivanova et al. (2013); Postnov and Yungelson (2014); note that the reverse scenario has been proposed as well (see e.g., Ref. Broekgaarden et al. (2022)). Another interesting property of asymmetric mass ratio and high total mass systems is that subdominant ringdown modes have a higher chance of being observable in such systems, see e.g. Capano et al. (2021).
@@ V Discussion
V.1 Dependence on the choice of astrophysical prior
@@
To facilitate further analysis and comparisons between different astrophysical priors, we make our posterior samples and other relevant information public at https://github.com/JayWadekar/GW_higher_harmonics_search. These can be used to reweight our results with other choices of priors (see e.g., Payne et al. (2019)). The biggest difference in our astrophysical prior choice compared to other catalogs Abbott et al. (2021); Nitz et al. (2021) is for the BH spins. We use a prior that is flat in the effective spin parameter . Oue motivation lies in the fact that is often the best-measured spin parameter, and hence we want to give equal weights to all physically possible values of that parameter Zackay et al. (2019). On the other hand, using an isotropic spin prior will significantly down-weight large (positive or negative) effective spins and can make it harder to reweight the results to a different prior. We also provide PE samples for posteriors sampled under the isotropic spin prior in the above repository. Another difference in our priors is for the case of total mass and mass ratios.
@@
It is also worth discussing how our other priors compare with the results of Abbott et al. (2023). We used a search prior uniform in , whereas, Ref. Abbott et al. (2023) modeled the distribution as a power law and found . We used a simple power law prior for the detector frame total masses and did not include any separate prior over redshift. Ref. Abbott et al. (2023) modeled the source primary mass distribution as a power law with supplemented by a Gaussian peak at . Ref. Abbott et al. (2023) modeled the merger rate as and found . The motivation for our prior choices was again to make it easy to reweight our results (both and overall IFAR values) later in a hierarchical population analysis (especially for the high mass and low mass ratio parameter space), see e.g., Roulet et al. (2020).
@@
A full population analysis not only requires the PE samples but also the of each event Roulet et al. (2020). For this reason, we also provide a file, IAS_HM_O3_triggers.hdf, in the repository that contains the full list of triggers (including the sub-threshold ones). Just for reference, we report subthreshold events () in Table 7 in the Appendix.
V.2 What is the gain/loss in sensitivity due to including HM in the search?
Unfortunately, we cannot currently answer the above question by comparing the results from our search to those from the previous IAS -only searches Olsen et al. (2022); Mehta et al. (2023). This is because our search not only includes the effect of HM, but also includes the improvements (for all modes) due to downweighting noise transients (see Section II.3 and Wadekar et al. (2023b)). Furthermore, compared to earlier, we use a different astrophysical prior (see section II.4) and use a different waveform approximant for modeling the waveforms (IMRPhenomXAS Pratten et al. (2020) instead of IMRPhenomD Khan et al. (2016)). For us to isolate the effect of HM, we will need to rerun our search (ideally on injections) with all the settings to be the same except having only modes in the template banks. We leave this study to a future work.
Previous studies have noted that the larger variety in waveform morphologies (brought by the inclusion of higher modes) can lead to a loss of sensitivity in the search, as the number of background triggers is increased, see e.g., Capano et al. (2014); Chandra et al. (2022). This can be understood from the fact that, with higher modes, the shape of the waveform also depends on the inclination and orbital phase, and these additional degrees of freedom increase the trials factor when finding the best-matching template. In our search, we do not blankly add additional degrees of freedom to the templates throughout the parameter space. Instead, our new detection statistic (see Section II.2 and Wadekar et al. (2023b)) includes HM while simultaneously marginalizing the likelihood over the extrinsic parameters (which are the same as in the -only search). This marginalization naturally downweights fine-tuned solutions, approaching the optimal statistic and preventing an undue increase of the trials factor (e.g., the edge-on solutions get naturally downweighted when we marginalize over distance, as their observable volume is smaller).
In our ranking statistic for the triggers, we also marginalize over the ratios of the SNR amplitudes of HM: and . In this integral, we use a prior distribution for derived from our astrophysical prior. Note that the prior on can be very different for templates in different regions of parameter space. For example, the median of our prior for is 0.04 for Bank 0 (a low-mass bank), while it is 0.3 for Bank 14 (a high-mass bank). If we did not include such a prior, we would indeed expect to encounter a large loss of sensitivity on including HM. Comparing the performance of our search on events detected in previous -only catalogs (see Tables 3 and 6), we can confirm that our search sensitivity for these events has not been significantly degraded and is roughly comparable.

### v2

@@ III.3 Comparison to the LVK and OGC -only catalogs
In Table LABEL:tab:lvc_events, we compare our pipeline’s results for the O3 Hanford–Livingston coincident events published in the GWTC-3 catalog444https://gwosc.org/GWTC-3/ Abbott et al. (2021). The GWTC-3 catalog combines results from four pipelines: three that search using templates which only contain the mode: GstLAL, Multi-Band Template Analysis (MBTA) and PyCBC Sachdev et al. (2019); Usman et al. (2016); Andres et al. (2022); and one that searches for transient signals with minimal assumptions about sources, cWB Klimenko et al. (2016). We also include the comparison with the 4-OGC catalog taken from the summary data in the repository555https://github.com/gwastro/4-ogc/blob/master/search/4OGC_top.txt listed in Nitz et al. (2021). Note that the quoted GWTC-3 squared SNR, , is from LVK’s PE runs for the events. The search SNR2 is generally lower, and varies between the different pipelines that the LVK uses.
@@
We broadly recover the confident LVK detections, except some vetoed/missed events which are discussed in more detail in the subsections below. Overall, three GWTC-3 events in Table 1 were vetoed in our pipeline. GW190701_203306 was vetoed due to excess power present after subtracting the best-fit waveform in Livingston (note that the data around this event was treated for scattered light glitches Abbott et al. (2021)). GW200308_173609 has GWTC-3 and it fell below our SNR collection threshold. Six GWTC-3 events were detected with in our pipeline (GW190403_051519, GW190917_114630, GW191113_071753, GW200202_154313, GW200208_222617, GW200322_091133). Note that all the GWTC-3 events we missed (except GW200202_154313) were confidently detected by only one of the LVK pipelines. We find GW200202_154313 gets penalized in the multi-detector coherence module of our pipeline possibly due to having a large SNR difference in the two detectors.
@@
At the end of Table LABEL:tab:lvc_events, we include comparison with the list of 7 events in 4-OGC which are not present in GWTC-3 catalog (from which we recover 5 with ). GW190909_114149 is a marginal candidate in GWTC-3 which has from our pipeline. We still include it in Table LABEL:tab:lvc_events as it was first a significant candidate in GWTC-2 Abbott et al. (2021), but later updated to be in their marginal list The LIGO Scientific Collaboration et al. (2021). We also checked the overlap of our new candidate event list with the sub-threshold candidate list (i.e., which have but FAR day-1) released by the LVK collaboration: GWTC-2.1 and GWTC-3 The LIGO Scientific Collaboration et al. (2021); Abbott et al. (2021); we show the corresponding comparison in Table 5 in the Appendix. In the same table, we also show the overlap of our new candidate list with sub-threshold candidates from the IAS -only pipeline Olsen et al. (2022); Mehta et al. (2023).
@@
As we discussed earlier in Section II.2, we set a particular SNR threshold () to collect triggers from individual detectors. We do this in order to regulate the number of single-detector triggers that are passed to the more costly coincidence step of the pipeline (see section II.2.1 for a detailed discussion). Due to this, we missed low-SNR events that were reported by -only pipelines in the previous literature. We label the events which fell below this threshold as “Below ” in the table.
@@
We do not include the following events in Table LABEL:tab:lvc_events because they were detected in Livingston–Virgo or Hanford–Virgo coincidence, or single detector search, all of which we did not run: GW190620_030421, GW190630_185205, GW190708_232457, GW190814, GW190910_112807, and GW190925_232845. We also only searched for BBH mergers in this paper, and did not recover the NSBH candidates GW191219_163120 and GW200115_042309 (in addition, the posterior probability for GW200210_092254 peaks in the NSBH region Abbott et al. (2021), which could lead to it being vetoed in our search). We hope to perform coincident searches which would also include Virgo (and KAGRA) in the future. We also intend to perform a separate search for NSBH mergers including our HM pipeline in a future paper.
@@ IV Interesting astrophysical properties of the new candidate events
In this section, we discuss interesting astrophysical properties of the 11 new candidate events, whose parameters we have reported in Table 1 and in Figs. 1, 2 and 3.
@@ IV.1 Masses in the IMBH range and pair instability mass gap
BHs roughly within the mass range – are designated as intermediate-mass black holes (IMBHs). There are a number of channels which can produce BHs in this mass regime, see e.g., Coleman Miller and Hamilton (2002); Franciolini et al. (2024). Several IMBH candidates have been suggested by electromagnetic observations, but these lack conclusive confirmation Greene et al. (2020). Gravitational waves have the potential to definitively detect objects in this regime, e.g., Abbott et al. (2020); Fragione et al. (2018). In our search, at % credibility, 4 candidate events have primary black hole masses in the IMBH range (# 2, 4, 5, 9 in Table 1).
Another interesting regime in BH masses is the pair instability mass gap. This is the regime in which stars undergo pair-instability supernova (where they are completely disrupted and no remnant is left behind). Otherwise, the stars undergo pulsational pair-instability supernova and iteratively lose their mass before collapsing to a black hole. These processes are expected to produce a gap roughly between – in the distribution of BH mass. Many of our new candidate events have support for masses in the gap. Note however that the exact boundaries of such a gap are not well determined, as they depend on multiple uncertain factors such as the rate of carbon to oxygen burning in the core , angular momentum transport, metallicity, wind loss, etc. Marchant and Moriya (2020); Farmer et al. (2019); van Son et al. (2020); Mehta et al. (2022); Hendriks et al. (2023); Golomb et al. (2023); Franciolini et al. (2024).
Recent analyses have not found any conclusive evidence for the upper mass gap in the observed BH mass distribution, e.g., Abbott et al. (2023); Roulet et al. (2021). This either challenges the current theoretical models of stellar evolution, or the gap might be filled by black holes formed via alternative mechanisms such as hierarchical mergers Anagnostou et al. (2020); Fragione et al. (2020); Veske et al. (2021); Fragione et al. (2021), stellar collisions Kremer et al. (2020), or from the gaseous disk channel Yang et al. (2019); McKernan et al. (2020); Tagawa et al. (2020); Ishibashi, W. and Gröbner, M. (2020); Secunda et al. (2020); Tagawa et al. (2021); Fabj et al. (2020); Tagawa et al. (2020); Samsing et al. (2022). Our new candidate events could be helpful in estimating the abundance of BH (or rate of BH mergers) in/beyond the pair-instability mass regime. This can also shed light on BH progenitor channels. Other potential explanation for high-mass merger detections could be that they are lower-mass mergers but are lensed by a foreground galaxy Dai et al. (2017).
@@ IV.2 High redshifts
We find two new candidates with redshift at % CL (# 4, 6 in Table 1). 5 of our new candidates have median redshift . Finding high-redshift BHs can be important to answer questions such as: How does the merger rate of BHs evolve with redshift Fishbach et al. (2018)? Does the merger rate change in a different manner compared to the the star formation rate (which peaks at Madau and Dickinson (2014))? Furthermore, different BH formation channels have different predictions for the merger rate at high redshift Mapelli et al. (2022); van Son et al. (2022); Boesky et al. (2024a, b). For example, the chemically homogeneous evolution channel or the stable mass transfer channel produces BBHs of relatively higher masses and lower redshifts than the common envelope evolution channel Mandel and de Mink (2016); van Son et al. (2022). In the dynamical channel, massive BHs can have relatively longer delay times compared to star formation Mapelli et al. (2019). The redshifts of binaries can be also be correlated with other properties of BHs, e.g., spins Biscoveanu et al. (2022); Kushnir et al. (2016); Zaldarriaga et al. (2018). We aim to test some of these models in a future population inference study.
IV.3 Preference for non-zero spins
Measuring spins of the black holes in a binary is one of the best ways to distinguish between isolated and dynamical formation channels. In the isolated channel, accretion can lead to spins of BHs which are aligned in the direction of the binary angular momentum (leading to positive values of the effective spin parameter ). Dynamical channels, on the other hand, predict the astrophysical distribution to be symmetric around 0. This is because the spins BH captured into a binary are not expected to be preferentially aligned along any particular direction. From the PE runs under our fiducial prior (flat in ), two new candidates have positive effective spin at CL (GW191213_204950, GW190530_030659). On the other hand, GW200114_020818 has at CL. Even under the isotropic spin prior, some of the new candidate events still have mild preference for being non-zero.
@@ IV.4 Asymmetric mass ratio systems
5 candidates have median mass ratio . The posteriors are broad in most cases, but there are two candidates with a strong preference for asymmetric mass-ratios (GW200114_020818 and GW200210_035448). Finding sources with asymmetric mass ratios can shed light on the formation channels Abbott et al. (2020b). The globular cluster (dynamical) channel predicts that most merging BBHs have nearly equal masses Rodriguez et al. (2016). This is because higher mass BHs in a globular cluster sink to their cores due to dynamical friction and vice versa for lighter BHs. Although there is a caveat here that if clusters can retain merger products, they can produce “second-generation” mergers whose mass ratios can deviate significantly from unity Rodriguez et al. (2019). In the isolated binary channel, in certain cases, an initially asymmetric mass ratio system can evolve toward a more symmetric configuration due to mass accretion episodes Ivanova et al. (2013); Postnov and Yungelson (2014); note that the reverse scenario has been proposed as well (see e.g., Gerosa et al. (2013); Broekgaarden et al. (2022)). Another interesting property of asymmetric mass ratio and high total mass systems is that subdominant ringdown modes have a higher chance of being observable in such systems, see e.g. Capano et al. (2021).
@@ V Discussion
V.1 Injection campaigns to quantify gain in sensitivity due to including HM
We quantify the improvement in the sensitive volume-time () of the search pipeline in our companion paper: Mehta et al. (2025). We use the injection catalog released by the LVK collaboration on Zenodo LIGO Scientific Collaboration (2021). We also compare the results of our IAS-HM pipeline with its (2,2)-only limit. We find significant improvement in for systems with asymmetric mass ratios and larger total masses, with gains in sensitivity even exceeding 100% at certain high masses (see Fig. 2 of Mehta et al. (2025)). We also do not find a loss of sensitivity in the IAS-HM pipeline for near equal mass-ratio mergers. We also find that the of the IAS pipeline is either larger than or comparable to that of the individual LVK pipelines (see Fig. 1 of Mehta et al. (2025)).
It is quite encouraging to see that the parameter space where we get a large improvement () over the LVK and the (2,2)-only IAS pipelines is similar to the parameter space where we detect most of the new candidate events, thus adding more credibility to our newly reported detections. We refer the reader to Mehta et al. (2025) for further details on the injection campaigns.
Previous studies have noted that the larger variety in waveform morphologies (brought by the inclusion of higher modes) can lead to a loss of sensitivity in the search, as the number of background triggers is increased, see e.g., Capano et al. (2014); Chandra et al. (2022). This can be understood from the fact that, with HMs, the shape of the waveform also depends on the inclination and orbital phase, and these additional degrees of freedom increase the trials factor when finding the best-matching template. As discussed in section II.2 and W24, we do not naively add additional degrees of freedom to the templates throughout the parameter space. Instead, our new detection statistic (see Section II.2 and W24) includes HM while simultaneously marginalizing the likelihood over the extrinsic parameters and mass ratio. This marginalization naturally downweights fine-tuned solutions, approaching the optimal statistic and preventing an undue increase of the trials factor (e.g., the edge-on solutions get naturally downweighted when we marginalize over distance, as their observable volume is smaller).
The downside of using a prior distribution on is that it is partly dependent on our choice of astrophysical prior, which is currently not precisely known. Especially, the mass ratio distribution is not very well constrained and affects our prior for . Using a different distribution for than the true astrophysical distribution can indeed lead to some loss of sensitivity. We chose a broad prior (flat in ) in our search, with the motivation that such a prior can be easily reweighted later in a hierarchical population analysis.
Even though including HM can cause some loss in sensitivity for near-threshold face-on equal mass systems, note that detecting even a few events with asymmetric-mass ratios can inform our population estimates much more than detecting a slightly larger number of equal-mass systems. It is worth drawing an analogy here to the case of spin distribution of templates used in -only searches. Neglecting templates with negative indeed improves the search sensitivity to near-threshold systems, but finding even a few systems can greatly help us distinguish between progenitor channels of BHs. We thus argue the goal of a search should not solely be to maximize the total number of detections, but to holistically cover most of the interesting regions of the binary parameter space. A similar argument also holds for when including spin-precession or eccentricity effects in the search templates. Detecting even a few events with significantly misaligned spins can shed more light on the formation channels of binaries as compared to detecting a larger number of near-threshold aligned/zero spin mergers. Some of the challenges for including these effects are similar to the ones for HM: significant increase in the template bank size and loss of sensitivity due to collecting additional background McIsaac et al. (2023); Schmidt et al. (2024a, b); Dhurkunde and Nitz (2022, 2023); Nitz et al. (2020). We will develop methods similar to the ones implemented in our search to help tackle those searches in future work Zhou et al. ; Islam et al. .
V.2 Dependence on the choice of astrophysical prior
@@
To facilitate further analysis and comparisons between different astrophysical priors, we make our posterior samples, the selection function of our pipeline and other relevant information public at https://github.com/JayWadekar/GW_higher_harmonics_search. These can be used to reweight our results with other choices of priors (see e.g., Payne et al. (2019)). The biggest difference in our astrophysical prior choice compared to other catalogs Abbott et al. (2021); Nitz et al. (2021) is for the BH spins. We use a prior that is flat in the effective spin parameter . Oue motivation lies in the fact that is often the best-measured spin parameter, and hence we want to give equal weights to all physically possible values of that parameter Zackay et al. (2019). On the other hand, using an isotropic spin prior will significantly down-weight large (positive or negative) effective spins and can make it harder to reweight the results to a different prior. We also provide PE samples for posteriors sampled under the isotropic spin prior in the above repository.
@@
Another difference in our priors is for the case of total mass and mass ratios. We used a search prior uniform in , whereas, Ref. Abbott et al. (2023) modeled the distribution as a power law and found . We used a simple power law prior for the detector frame total masses and did not include any separate prior over redshift. Ref. Abbott et al. (2023) modeled the source primary mass distribution as a power law with supplemented by a Gaussian peak at . Ref. Abbott et al. (2023) modeled the merger rate as and found . The motivation for our prior choices was again to make it easy to reweight our results (both and overall IFAR values) later in a hierarchical population analysis (especially for the high mass and low mass ratio parameter space), see e.g., Roulet et al. (2020).
@@
A full population analysis not only requires the PE samples but also the of each event Roulet et al. (2020); Gaebel et al. (2019). For this reason, we also provide a file, IAS_HM_O3_triggers.json, in the repository that contains the full list of triggers (including the sub-threshold ones). We also include the subthreshold events () in Table 7 in the Appendix for reference.
V.3 Population study
To probe interesting population inference questions in section IV, one needs to perform a hierarchical bayesian analysis and account for selection function of our search. We leave this analysis to an upcoming work Mehta et al. . The new candidate events reported here have moderate IFARs (the best event has yr), thus a methodology to combine confident and marginal events in population inference will be required Roulet et al. (2020). Our sensitivity is higher than other searches especially for high mass and/or asymmetric mass ratio events due to addition of HM. Having higher sensitivity can improve over some regions of the population hyperparameter space even if we have null/marginal detections.